OpenStudy (anonymous):

how many real number solutions does the equation -8x^2-5x-1=0 have?

jimthompson5910 (jim_thompson5910):

D = b^2 - 4ac D = (-5)^2 - 4(-8)(-1) D = 25 - 32 D = -7 The discriminant is negative, so there are no real solutions

jimthompson5910 (jim_thompson5910):

There are only imaginary solutions